# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

### v25

@@ 3.30 Tokenizing
The study by Petrov et al. [729] compared the performance 17 off-the-shelf language tokenizers on a large number of languages including Sinhala. They have found that some languages such as English have shorter tokenizing lengths (close to $1.0$) across the board while some can raise to undesirable heights (they observe $15.0$ as the highest). They report undesirable numbers for Sinhala, mostly in the $8.83$ to $12.86$ range. They report reasonable results on MBart50 [730] and mT5 [349, 730] with $1.35$ and $1.66$ respectively. The best result ($1.00$), is given by CANINE [731]. The work by Velayuthan and Sarveswaran [732] showed that the algorithms used in pre-tokenization are more important than the tokenization algorithm itself in the context of LLMs for languages such as Sinhala. Ranasinghe [163] claims to have created a comprehensive tokenizer for Sinhala and have made it publicly available174174174https://github.com/Buddhilive/sltk. However, the tokenizer is heavily dependent on a manually created dictionary to handle compound words. While the authors have made the dictionary publicly available175175175https://github.com/Buddhilive/sltk/blob/main/sltkpy/shared/dict.json, it only contains seven entries. So while it may be adequate to handle Sri Lanka-centric news title data set176176176https://github.com/Buddhilive/sltk/blob/main/data/sin.csv provided with the code, it may not work well in the general domain.
3.31 Detection or Management of Mental Illnesses
Using classical machine learning algorithms on TF-IDF features, Fazal and Farook [698] predicted depression in Sinhala-English code-mixed data from Twitter and Facebook. Hewapathirana and Sumanathilaka [733] introduced EmoScan, a method to detect depression in Romanized Sinhala (Singlish) tweets. The study by Herath and Wijayasiriwardhane [734] used classical machine learning models to predict suicidal ideation from Sinhala Facebook posts. Fazal and Farook [735] compared multiple algorithms to detect depression in Sinhala English code mixed social media posts and reported that ExtraTreesClassifier [736] obtains the best results. Welarathna et al. [428] used CNNs to classify emotions of Sinhala speech by Autistic children. They handle the following emotions: sad, disgust, surprise, neutral, happy, calm, fear, and angry. Jayawardena et al. [737] used a NAO Robot [738] to create an interactive system with Sinhala language content detection (Speech to Text), dialogue management, and voice synthesis (Text-to-Speech) for Autism Intervention in children. The system by Kulasooriya et al. [739] used SVM and simple ANN to detect psychological stress in Sinhala speech.
@@ 3.33 Question Answering Systems (Q&A)
Romero et al. [161] introduced CVQA, a multilingual visual Question Answering benchmark which includes Sinhala. They have publicly released both their data177177177https://huggingface.co/datasets/afaji/cvqa and leader board178178178https://eval.ai/web/challenges/challenge-page/2305/overview. Vayani et al. [164] created a dataset ALM-Bench179179179https://huggingface.co/datasets/MBZUAI/ALM-Bench by machine translating English QA pairs from LLaVA-Bench (In-the-Wild) dataset [165] and extended it by searching for images under country name, language name, and cultural category on the web. They then reported that for Sinhala Q&A (especially visual Q&A), GPT-4o [166] is superior to other tested LLMs. Wansekara and Jayasekara [751] developed a visual Sinhala question and answer system using YOLO for object identification and a BERT model trained on a Sinhala corpus for Sinhala text. Dhanawardhana et al. [752] frame their work just as a multi-model deep learning approach. But in actuality, their lost-and-found application is utilising techniques that are relevant to Sinhala VQA. The system created by Dasanayaka et al. [634] is claimed to be capable of first, generating radiology reports for orthopantomography (related to dentistry) using Llama on captions generated by a Blip-2-based caption generator on dental panoramic tomography (DPT) images and then serve a Sinhala chatbot where the patients may ask questions on the said report and get answers.
@@ 3.34 Language Identification (LID) Task
Burchell et al. [146] show that they obtain better results for the LID task for most of the languages compared to NLLB [134]. However, for Sinhala the results by Burchell et al. [146] are the same as that of NLLB. Dunn and Edwards-Brown [8] experiment with improving the LID task with the help of a geographical model. However, they conclude that adding the geographical model does not help in the Sinhala LID task given that the LID model and the geographical model already agree with each other 99%. Kargaran et al. [753] introduces a tool named GlotScript which is used for written LangID it achieves above 80% accuracy for Sinhala.
@@ 3.35 Miscellaneous Applications
Kumari and Hettiarachchi [760] proposed an algorithm for Sinhala topic modelling based on LDA [264] and RAKE [761]Batawalaarachchi [762] proposed two methods using statistical features to select words to be included in Sinhala document titles. The study by pal [763] conducted a comparative analysis of LDA, LSA, and NMF for Sinhala and concluded that NMF archives best results. Arambewela et al. [764] proposed a Sinhala writing assistant tool utilising CNNS. The study by Bandara et al. [765] has introduced a Sinhala virtual assistant with rule-based decision-making for a pre-set collection of user queries and a general Transformer-based model for general user questions.  Jayaweera et al. [766] used classical machine learning methods to propose dynamic stop word removal. They claim to have released a corpus of 100,000+ Sinhala documents in their paper. But they provide no information on where to obtain this corpus. Liyanage et al. [767] have created a small Sinhala treebank of 100 sentences. Minixhofer et al. [768] have used Sinhala from OPUS100 [131] dataset to pre-train their multilingual punctuation-agnostic sentence segmentation model. Given that Sinhala is not one of their main target languages, they do not report sentence segmentation results for Sinhala in much detail. However, they do report that their method improves Sinhala sentence splitting F1 score to 86.0 above their SpaCy baseline of 75.8. A meta-study on how language data from social media are used in research was conducted by Hewapathirana [769]. A Sinhala programming language named Helaa based on Java was proposed by Yasasri and Karunarathna [770]. Senarathne [771] introduced a model to translate Sinhala pseudocode to computer code written in C#. The work by De Silva and Athukorala [772] claims to be able to provide Sinhala language explanations to JAVA code as well as generate JAVA code snippets from Sinhala descriptions.
The study by [229] performs best in this regard. used NER to cluster Sinhala news articles. Subasinghe and Samarawickrama [773] analysed the Noto Sans Sinhala180180180https://fonts.google.com/noto/specimen/Noto+Sans+Sinhala font in the context of legibility on small-scale digital device screens. Zhao et al. [774, 775] discussed a Sinhala LLM-based system they utilized to conduct participatory research in Sri Lanka. While they discussed how the LLM system helped their goals and briefly touched on how they utilised a hybrid translation approach of NMT and rule-based methods to circumvent the lack of training data, the specifics of their LLM system are not given.
@@ 4 Primary Sources
A number of NLP applications have utilized first sources intended to teach children [790, 791, 792, 793, 794] or foreigners [795, 221, 796]181181181Note that [796] is an extension of [221].. This makes sense given that an introduction written for children would start with basic principles and thus be ideal for crafting rule based NLP systems and an introduction written for foreigners would have Sinhala language described in terms of English, making easy the process of rule based translation of English NLP tools to Sinhala.
@@ Appendix C Author Meta Analysis
Figure 10: The Probability of research from an institution citing that of other institutions. Note that the calculations are limited by 4 factors: (1) The availability of free to download pdf of the paper, (2) The aforementioned pdf containing references list (Some extended abstracts do not come with the references list), (3) The text extraction capabilities of pdftotext183183183https://www.xpdfreader.com/pdftotext-man.html, (4) The accuracy of the research paper title look up. With all those limitations in mind, we still can make a few Interesting observations. The institute with the highest number of publications, UCSC - University of Colombo seems to be getting the most citations from most sources. However, they themselves seem to almost exclusively cite their own papers (0.7543). The only exception is the smaller number of works they cite (0.2114) from the Department of CSE - University of Moratuwa, the institute with the second highest number of publications. This leaves them with a 0.0343 probability of citing anyone else. Comparatively, Department of CSE - University of Moratuwa seems to be more egalitarian in citing. They have a lower self-citation probability (0.5251) and a higher probability of citing UCSC - University of Colombo (0.3513). This results in a probability of citing others at 0.1236. Both the Johns Hopkins University and University of Edinburgh prefer to cite Google instead of work from Sri Lanka. Faculty of IT - University of Moratuwa prefers to cite UCSC - University of Colombo (0.2339) rather than Department of CSE - University of Moratuwa (0.0968) which is from the same parent institution.

### v26

@@ 3.30 Tokenizing
The study by Petrov et al. [794] compared the performance 17 off-the-shelf language tokenizers on a large number of languages including Sinhala. They have found that some languages such as English have shorter tokenizing lengths (close to $1.0$) across the board while some can raise to undesirable heights (they observe $15.0$ as the highest). They report undesirable numbers for Sinhala, mostly in the $8.83$ to $12.86$ range. They report reasonable results on MBart50 [795] and mT5 [366, 795] with $1.35$ and $1.66$ respectively. The best result ($1.00$), is given by CANINE [796]. The work by Velayuthan and Sarveswaran [797] showed that the algorithms used in pre-tokenization are more important than the tokenization algorithm itself in the context of LLMs for languages such as Sinhala. Ranasinghe [166] claims to have created a comprehensive tokenizer for Sinhala and have made it publicly available223223223https://github.com/Buddhilive/sltk. However, the tokenizer is heavily dependent on a manually created dictionary to handle compound words. While the authors have made the dictionary publicly available224224224https://github.com/Buddhilive/sltk/blob/main/sltkpy/shared/dict.json, it only contains seven entries. So while it may be adequate to handle Sri Lanka-centric news title data set225225225https://github.com/Buddhilive/sltk/blob/main/data/sin.csv provided with the code, it may not work well in the general domain.
3.31 Detection or Management of Illnesses
Using classical machine learning algorithms on TF-IDF features, Fazal and Farook [745] predicted depression in Sinhala-English code-mixed data from Twitter and Facebook. Hewapathirana and Sumanathilaka [798] introduced EmoScan, a method to detect depression in Romanized Sinhala (Singlish) tweets. The study by Herath and Wijayasiriwardhane [799] used classical machine learning models to predict suicidal ideation from Sinhala Facebook posts. Fazal and Farook [800] compared multiple algorithms to detect depression in Sinhala English code mixed social media posts and reported that ExtraTreesClassifier [801] obtains the best results.
Welarathna et al. [454] used CNNs to classify emotions of Sinhala speech by Autistic children. They handle the following emotions: sad, disgust, surprise, neutral, happy, calm, fear, and angry. Jayawardena et al. [802] used a NAO Robot [803] to create an interactive system with Sinhala language content detection (Speech to Text), dialogue management, and voice synthesis (Text-to-Speech) for Autism Intervention in children. Continuing their work, Jayawardena et al. [804] used the RASA framework [668] to build a dialogue system that provides speech therapy to Sinhala-speaking Children with Autism Spectrum Disorder (ASD). Further, Manukalpa et al. [805] extended this work with cultural alignment.
The system by Kulasooriya et al. [806] used SVM and simple ANN to detect psychological stress in Sinhala speech. Rathnayake et al. [807] used voice classification by biomarkers to detect Parkinson’s disease with classical machine learning algorithms.
@@ 3.33 Question Answering Systems (Q&A)
Romero et al. [164] introduced CVQA, a multilingual visual Question Answering benchmark which includes Sinhala. They have publicly released both their data226226226https://huggingface.co/datasets/afaji/cvqa and leader board227227227https://eval.ai/web/challenges/challenge-page/2305/overview. Vayani et al. [167] created a dataset ALM-Bench228228228https://huggingface.co/datasets/MBZUAI/ALM-Bench by machine translating English QA pairs from LLaVA-Bench (In-the-Wild) dataset [168] and extended it by searching for images under country name, language name, and cultural category on the web. They then reported that for Sinhala Q&A (especially visual Q&A), GPT-4o [169] is superior to other tested LLMs. Wansekara and Jayasekara [819] developed a visual Sinhala question and answer system using YOLO for object identification and a BERT model trained on a Sinhala corpus for Sinhala text. Dhanawardhana et al. [820] frame their work just as a multi-model deep learning approach. But in actuality, their lost-and-found application is utilising techniques that are relevant to Sinhala VQA. The system created by Dasanayaka et al. [676] is claimed to be capable of first, generating radiology reports for orthopantomography (related to dentistry) using Llama on captions generated by a Blip-2-based caption generator on dental panoramic tomography (DPT) images and then serve a Sinhala chatbot where the patients may ask questions on the said report and get answers. Shafique et al. [785] introduced ViMUL-Bench a multimodal video benchmark for LLMs that includes Sinhala. With 49.6, they show that GPT4o has the best results for Sinhala; however, the Sinhala accuracy is the lowest of all the tested languages across almost all the LLMs.
@@ 3.34 Language Identification (LID) Task
Burchell et al. [149] show that they obtain better results for the LID task for most of the languages compared to NLLB [137]. However, for Sinhala the results by Burchell et al. [149] are the same as that of NLLB. Dunn and Edwards-Brown [8] experiment with improving the LID task with the help of a geographical model. However, they conclude that adding the geographical model does not help in the Sinhala LID task given that the LID model and the geographical model already agree with each other 99%. Kargaran et al. [821] introduces a tool named GlotScript which is used for written LangID it achieves above 80% accuracy for Sinhala. Van Der Goot [822] discussed the current open challenges in LangID task with 2000+ languages. However, due to them only having the Sinhala language written with Sinhala characters, all models in their experiment obtained 99%+ F1 vlaues for Sinhala with Textcat [823]229229229https://www.let.rug.nl/vannoord/TextCat/ and GLOT500 [112] obtaining 99.9%. It should be noted that in the original GLOT500230230230https://github.com/cisnlp/Glot500 paper, Imani et al. [112] only achieved 45.0% accuracy for Sinhala.
@@ 3.35 Miscellaneous Applications
Kumari and Hettiarachchi [830] proposed an algorithm for Sinhala topic modelling based on LDA [275] and RAKE [831]Batawalaarachchi [832] proposed two methods using statistical features to select words to be included in Sinhala document titles. The study by Pallawala and Haddela [833] conducted a comparative analysis of LDA, LSA, and NMF for Sinhala and concluded that NMF archives best results. Arambewela et al. [834] proposed a Sinhala writing assistant tool utilising CNNS. The study by Bandara et al. [835] has introduced a Sinhala virtual assistant with rule-based decision-making for a pre-set collection of user queries and a general Transformer-based model for general user questions.  Jayaweera et al. [836] used classical machine learning methods to propose dynamic stop word removal. They claim to have released a corpus of 100,000+ Sinhala documents in their paper. But they provide no information on where to obtain this corpus. Liyanage et al. [837] have created a small Sinhala treebank of 100 sentences. Minixhofer et al. [838] have used Sinhala from OPUS100 [134] dataset to pre-train their multilingual punctuation-agnostic sentence segmentation model. Given that Sinhala is not one of their main target languages, they do not report sentence segmentation results for Sinhala in much detail. However, they do report that their method improves Sinhala sentence splitting F1 score to 86.0 above their SpaCy baseline of 75.8. A meta-study on how language data from social media are used in research was conducted by Hewapathirana [839]. A Sinhala programming language named Helaa based on Java was proposed by Yasasri and Karunarathna [840]. Senarathne [841] introduced a model to translate Sinhala pseudocode to computer code written in C#. The work by De Silva and Athukorala [842] claims to be able to provide Sinhala language explanations to JAVA code as well as generate JAVA code snippets from Sinhala descriptions.
The study by [238] performs best in this regard. used NER to cluster Sinhala news articles. Subasinghe and Samarawickrama [843] analysed the Noto Sans Sinhala231231231https://fonts.google.com/noto/specimen/Noto+Sans+Sinhala font in the context of legibility on small-scale digital device screens. Zhao et al. [844, 845] discussed a Sinhala LLM-based system they utilized to conduct participatory research in Sri Lanka. While they discussed how the LLM system helped their goals and briefly touched on how they utilised a hybrid translation approach of NMT and rule-based methods to circumvent the lack of training data, the specifics of their LLM system are not given. Wijerathne et al. [846] introduced ScheduleMe an LLM multi-agent system to manage a Google calendar. They reported a success rate of 70% for Sinhala. It should be noted that among the languages tested by them, only Chinese achieved a worse result (65%) than Sinhala. Jayatilleke and de Silva [497] introduced SiDiaC232232232https://github.com/NeviduJ/SiDiaC, the first Sinhala diachronic corpus. It covers 58k tokens across 46 Sinhala literary works from the 5th to the 20th century CE under Religious, History, Poetry, Language, and Medical genres. The work by Jayasiri and Peellawalage [847] attempted to identify Sinhala song genres by classifying their Mel-frequency Cepstral Coefficients (MFCCs) [848] spectrogram images using CNN models such as VGG-19, InceptionV3, EfficientNet, and DenseNet.
@@ 4 Primary Sources
A number of NLP applications have utilized first sources intended to teach children [863, 864, 865, 866, 867] or foreigners [868, 230, 869]233233233Note that [869] is an extension of [230].. This makes sense given that an introduction written for children would start with basic principles and thus be ideal for crafting rule based NLP systems and an introduction written for foreigners would have Sinhala language described in terms of English, making easy the process of rule based translation of English NLP tools to Sinhala.
@@ Appendix C Author Meta Analysis
Figure 10: The Probability of research from an institution citing that of other institutions. Note that the calculations are limited by 4 factors: (1) The availability of free to download pdf of the paper, (2) The aforementioned pdf containing references list (Some extended abstracts do not come with the references list), (3) The text extraction capabilities of pdftotext235235235https://www.xpdfreader.com/pdftotext-man.html, (4) The accuracy of the research paper title look up. With all those limitations in mind, we still can make a few Interesting observations. The institute with the highest number of publications, UCSC - University of Colombo seems to be getting the most citations from most sources. However, they themselves seem to almost exclusively cite their own papers (0.7543). The only exception is the smaller number of works they cite (0.2114) from the Department of CSE - University of Moratuwa, the institute with the second highest number of publications. This leaves them with a 0.0343 probability of citing anyone else. Comparatively, Department of CSE - University of Moratuwa seems to be more egalitarian in citing. They have a lower self-citation probability (0.5251) and a higher probability of citing UCSC - University of Colombo (0.3513). This results in a probability of citing others at 0.1236. Both the Johns Hopkins University and University of Edinburgh prefer to cite Google instead of work from Sri Lanka. Faculty of IT - University of Moratuwa prefers to cite UCSC - University of Colombo (0.2339) rather than Department of CSE - University of Moratuwa (0.0968) which is from the same parent institution.